A body falls freely from the rest, Find:

A. Distance it falls in 5 sec.

B. Velocity after falling 5 km.

C. Time required to attain a velocity of 50 m/s.

D. Time in which it falls 200 m.


1
Expert's answer
2022-03-09T12:20:58-0500

a)


"h=0.5gt^2=0.5(9.8)(5)^2=122.5\ m"


b)


"v=\sqrt{2gh}=\sqrt{2(9.8)(5000)}=313\frac{m}{s}"


c)


"t=\frac{v}{g}=\frac{50}{9.8}=5.1\ s"


d)


"T=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(200)}{9.8}}=6.4\ s"